OpenStudy (anonymous):

Solve by factoring. 2x^2 + 9x =5

jagr2713 (jagr2713):

answer choice

jagr2713 (jagr2713):

x=−5,1/2

OpenStudy (anonymous):

can you show me how you got it

jagr2713 (jagr2713):

well he will tell you LOL

jagr2713 (jagr2713):

save me from typing alot

OpenStudy (triciaal):

first set the equation =0 2x^2 +9x -5 = 0 in the form Ax^2 + Bx + C =0 what are the factors of AC such that when you add you get B 2, 5, -1 (2x-1)(x + 5) =0 2x -1 =0 ; x + 5 = 0 x = 1/2 ; x = -5